# RMBRec: Robust Multi-Behavior Recommendation towards Target Behaviors

### v1

@@ 1. Introduction
Recommender systems (RS) have become essential tools of the Web ecosystem, helping users discover preferred items, news, or services from massive candidate pools (Chen et al., 2019; Covington et al., 2016; Sun et al., 2024). Traditional studies usually focus on a single type of feedback (e.g., clicks or purchases), a setting often referred to as single-behavior recommendation (Rendle et al., 2009; Chen et al., 2020; Koren et al., 2009). However, user interactions in real-world scenarios are inherently multi-dimensional, including view, collect, add-to-cart, and purchase behaviors (Jin et al., 2020; He et al., 2024b; Ma et al., 2025; Han et al., 2024). Leveraging such heterogeneous feedback, known as multi-behavior recommendation, can enrich supervision and improve the prediction of target behaviors (e.g., purchases) (Wu et al., 2024; Gao et al., 2019; Xu et al., 2023).
Despite this potential, most existing multi-behavior recommendation works rely on fusion-based paradigm, which learns behavior-specific embeddings and aggregates them through concatenation (Chen et al., 2020; He et al., 2020), MLPs (Han et al., 2024; Gao et al., 2019), or attention mechanisms (Jin et al., 2020; Meng et al., 2023; Ma et al., 2025). In practice, however, auxiliary interactions (e.g., view or cart) are often noisy or weakly correlated with the target behavior (Han et al., 2024; Cai et al., 2025; Wang et al., 2025; Zhao et al., 2025). Users may casually browse or add items to carts without genuine purchase intent, introducing inconsistent supervision signals that distort the learned preference space. As a result, multi-behavior recommenders tend to overfit spurious correlations rather than capture true purchasing intent, leading to poor robustness in real-world environments (Xu et al., 2025; Wang et al., 2024; O’Mahony et al., 2004). This vulnerability mainly arises from two types of inconsistencies, as shown in Figure 1. (1) Distributional shift: some users directly purchase without performing auxiliary actions, causing the distributions of auxiliary and target behaviors to diverge (Yan et al., 2025; Cai et al., 2025). (2) Auxiliary noise: many auxiliary interactions (e.g., random views or abandoned carts) are unrelated to real purchase intent (Zhang et al., 2023a; Zhao et al., 2025). These two factors jointly create semantic misalignment across behaviors, making naive fusion of heterogeneous signals unreliable.
@@
To quantify the severity of such misalignment, we introduce a simple diagnostic metric called the Behavioral Alignment Ratio (BAR), which measures the degree to which each auxiliary behavior aligns with the target one. BAR reveals significant disparities across datasets, indicating that auxiliary signals differ drastically in reliability. For example, in some datasets nearly all purchases follow auxiliary actions (high BAR), while in others the majority of purchases occur independently (low BAR). This observation highlights a key insight that a robust model must exploit well-aligned signals while resisting noisy or misaligned ones. From an information-theoretic perspective (Wu et al., 2020; Yang et al., 2021), robustness in multi-behavior learning requires maximizing the predictive information extracted from reliable behaviors while minimizing the influence of noisy or distributionally shifted ones.
@@ 2. Related Works
Multi-Behavior Recommendation. Multi-behavior recommendation (MBR) enhances user modeling by incorporating heterogeneous interactions such as view, cart, and purchase (Gao et al., 2019; Jin et al., 2020). Early studies extend matrix factorization and collaborative filtering by sharing latent factors across behaviors (Gao et al., 2019). With the success of graph neural networks (GNNs), models such as MBGCN (Jin et al., 2020), MB-CGCN (Cheng et al., 2023), and CRGCN (Yan et al., 2023) construct behavior-specific graphs and propagate user–item relations to capture high-order dependencies. Attention-based approaches (Yan et al., 2024; Meng et al., 2023) further introduce behavior-level attention to adaptively weigh the contribution of each behavior. While these methods effectively utilize multi-behavior signals, they usually assume that all auxiliary interactions are semantically consistent with the target behavior. However, in real-world data, auxiliary signals may be noisy or even misaligned (e.g., frequent views without purchases), which can introduce biased supervision and reduce model robustness (Gao et al., 2022; Li et al., 2023; He et al., 2024a).
@@ 3.2. Representation Robustness Module (RRM)
In multi-behavior recommendation, auxiliary behaviors often have noise in the semantic intent from the target behavior. For example, users may view or add-to-cart without genuine purchase intention, which induces semantic drift on the user side. In contrast, item characteristics are relatively stable across behaviors. To address this local misalignment, we introduce RRM to enhance local semantic consistency. It maximizes the predictive information between auxiliary and target user representations, i.e., $I(P_{u}^{(b)};P_{u}^{(t)})$, so that the learned embeddings retain target-relevant signals while suppressing behavior-specific noise. In particular, RRM realizes this objective via a target-anchored contrastive formulation.
@@ 3.4. Target Prediction and Optimization
We employ equal-weight fusion to aggregate behavior-specific embedding. This equal-weight fusion ensures that all behavior types contribute equally to the final representation, while the purification through RRM and ORM ensures that only semantically aligned and optimizational stable signals are preserved. Using these fused embeddings, we compute the final preference score for target behavior prediction:
(9) $s(u,i)=\mathbf{z}_{u}^{\top}\mathbf{z}_{i}.$
@@ 4.2. Overall Performance
  • Multi-behavior modeling consistently outperforms single-behavior methods. LightGCN and SimGCL, which rely solely on the target behavior, achieve noticeably lower accuracy compared to multi-behavior models across all datasets. This validates the effectiveness of incorporating auxiliary behavioral signals, which provide complementary supervision and alleviate target sparsity.
  • Stable and consistent performance across diverse datasets. The performance trends of RMBRec remain stable under different behavioral characteristics. On Taobao, where auxiliary behaviors are moderately noisy, RMBRec achieves substantial gains (up to +29.5% on NDCG@10). On TMall, where behaviors are weakly correlated with the target, it maintains competitive performance without degradation. Even on the clean and well-structured Beibei, it continues to outperform all baselines. These results highlight the strong generalization ability of RMBRec and its robustness to varying behavior quality and distributional patterns.
4.3. Ablation Study
@@ 4.4. Robustness Study
4.5. Invariance Regularization
Variant Taobao TMall Beibei
HR@10 NDCG@10 HR@10 NDCG@10 HR@10 NDCG@10
w/o ORM 0.1321 0.0887 0.1360 0.0766 0.0871 0.0461
IRM-V1 0.1456 0.0962 0.1411 0.0794 0.0913 0.0483
IRM-V2 0.1389 0.0938 0.1304 0.0729 0.0902 0.0478
IRM-REx 0.1462 0.0969 0.1549 0.0838 0.0924 0.0489
Invariance learning serves as a key component in RMBRec, enabling the model to capture global optimization stability that holds across heterogeneous behavioral environments. To examine its impact, we implement three widely used variants of Invariant Risk Minimization (IRM) (Arjovsky et al., 2019): IRM-V1 (Arjovsky et al., 2019), IRM-V2 (Ahuja et al., 2020), and IRM-REx (Krueger et al., 2021) (see Appendix C.2 for implementation details).
As shown in Table 4, all IRM variants substantially outperform the model without ORM, confirming that invariance regularization effectively suppresses behavior-specific noise and improves robustness. Among them, IRM-REx achieves the best overall performance across all datasets, benefiting from its variance-based penalty that enforces consistent risks across behavioral environments, thus promoting more stable global optimization. IRM-V1 also yields notable gains, though its gradient-based constraint can be harder to optimize in practice. By contrast, IRM-V2 performs weakest, as its fixed classifier ($w=1$) limits flexibility in learning invariant representations. These results highlight the importance of global invariance regularization in stabilizing learning across diverse behaviors, and demonstrate that well-designed risk-based objectives such as REx can further enhance robustness in multi-behavior recommendation.
@@ Acknowledgements.
This research is supported by the National Natural Science Foundation of China (No. 62272254, No. U25A20445, No. 72188101).
@@ A.2. Detailed Description of  RMBRec Algorithm
Algorithm 1 outlines the complete training procedure of  RMBRec, which systematically integrates multi-behavior learning with dual-level robustness enhancement. The algorithm operates through six key steps:
  1. (1)
    Learning Behavior-Specific Embedding : For each behavior type, LightGCN learns specialized user and item embeddings from the corresponding interaction graph, starting from the unified initialization.
  2. (2)
    Representation Robustness Module: Applies target-anchored contrastive learning to align auxiliary behavior embeddings with the target behavior, ensuring semantic consistency and filtering out noisy signals.
  3. (3)
    Optimization Robustness Module: Enforces invariance regularization across behavioral environments by penalizing variance in BPR losses, capturing stable preference patterns that generalize beyond behavior-specific noise.
  4. (4)
    Robustness-Enhanced Fusion: Averages the purified behavior-specific embeddings to obtain the final unified representations, ensuring that only semantically aligned signals contribute to target prediction.
  5. (5)
    Target Behavior Prediction: Computes the main recommendation loss using the fused embeddings and target behavior interactions as supervision.
  6. (6)
    Model Optimization: Combines all objectives using hyperparameters $\lambda_{1}$ and $\lambda_{2}$ to balance the contributions of representation robustness and optimization robustness.

### v2

@@ 1. Introduction
Recommender systems (RS) have become essential tools of the Web ecosystem (Zhuo et al., 2024; Xu et al., 2025), helping users discover preferred items, news, or services from massive candidate pools (Chen et al., 2019; Covington et al., 2016). Traditional studies usually focus on a single type of feedback (e.g., clicks or purchases), a setting often referred to as single-behavior recommendation (Rendle et al., 2009; Chen et al., 2020; Koren et al., 2009). However, user interactions in real-world scenarios are inherently multi-dimensional, including view, collect, add-to-cart, and purchase behaviors (Jin et al., 2020; He et al., 2024b; Ma et al., 2025; Han et al., 2024). Leveraging such heterogeneous feedback, known as multi-behavior recommendation, can enrich supervision and improve the prediction of target behaviors (e.g., purchases) (Wu et al., 2024).
Despite this potential, most existing multi-behavior recommendation works rely on fusion-based paradigm, which learns behavior-specific embeddings and aggregates them through concatenation (Chen et al., 2020; He et al., 2020), MLPs (Han et al., 2024; Gao et al., 2019), or attention mechanisms (Jin et al., 2020; Meng et al., 2023; Ma et al., 2025). In practice, however, auxiliary interactions (e.g., view or cart) are often noisy or weakly correlated with the target behavior (Han et al., 2024; Cai et al., 2025; Wang et al., 2025; Zhao et al., 2025). Users may casually browse or add items to carts without genuine purchase intent, introducing inconsistent supervision signals that distort the learned preference space. As a result, multi-behavior recommenders tend to overfit spurious correlations rather than capture true purchasing intent, leading to poor robustness in real-world environments (Wang et al., 2024; O’Mahony et al., 2004). This vulnerability mainly arises from two types of inconsistencies, as shown in Figure 1. (1) Distributional shift: some users directly purchase without performing auxiliary actions, causing the distributions of auxiliary and target behaviors to diverge (Yan et al., 2025; Cai et al., 2025). (2) Auxiliary noise: many auxiliary interactions (e.g., random views or abandoned carts) are unrelated to real purchase intent (Zhang et al., 2023a; Zhao et al., 2025). These two factors jointly create semantic misalignment across behaviors, making naive fusion of heterogeneous signals unreliable.
@@
To quantify the severity of such misalignment, we introduce a diagnostic metric called the Behavioral Alignment Ratio (BAR), which measures the degree to which each auxiliary behavior aligns with the target one. BAR reveals significant disparities across datasets, indicating that auxiliary signals differ in reliability. For example, in some datasets nearly all purchases follow auxiliary actions (high BAR), while in others the majority of purchases occur independently (low BAR). This observation highlights an insight that a robust model must exploit well-aligned signals while resisting noisy or misaligned ones. From an information-theoretic perspective (Wu et al., 2020; Yang et al., 2021), robustness in multi-behavior learning requires maximizing the predictive information from reliable behaviors while minimizing the influence of noisy or distributionally shifted ones.
@@ 2. Related Works
Multi-Behavior Recommendation. Multi-behavior recommendation (MBR) enhances user modeling by incorporating heterogeneous interactions such as view, cart, and purchase (Gao et al., 2019; Jin et al., 2020). Early studies extend matrix factorization and collaborative filtering by sharing latent factors across behaviors (Gao et al., 2019). With the success of graph neural networks (GNNs) (Liu et al., 2025), models such as MBGCN (Jin et al., 2020), MB-CGCN (Cheng et al., 2023), and CRGCN (Yan et al., 2023) construct behavior-specific graphs and propagate user–item relations to capture high-order dependencies. Attention-based approaches (Yan et al., 2024; Meng et al., 2023) further introduce behavior-level attention to adaptively weigh the contribution of each behavior. While these methods effectively utilize multi-behavior signals, they usually assume that all auxiliary interactions are semantically consistent with the target behavior. However, in real-world data, auxiliary signals may be noisy or even misaligned (e.g., frequent views without purchases), which can introduce biased supervision and reduce model robustness (Gao et al., 2022; Li et al., 2023; He et al., 2024a; Liu et al., 2016).
@@ 3.2. Representation Robustness Module (RRM)
In multi-behavior recommendation, semantic drift is inherently user-centric. While item properties remain invariant across behaviors (e.g., an item’s intrinsic attributes stay consistent whether it is merely viewed or purchased), user intent varies significantly. For example, users may casually browse or add items to carts without genuine purchase intention, which induces stochastic noise and semantic drift specifically in user representations. To address this user-side misalignment, we introduce RRM to enhance local semantic consistency. It maximizes the predictive information between auxiliary and target user representations, i.e., $I(P_{u}^{(b)};P_{u}^{(t)})$, ensuring that the learned embeddings retain target-relevant signals while suppressing behavior-specific noise. In particular, RRM realizes this objective via a target-anchored contrastive formulation.
@@ 3.4. Target Prediction and Optimization
We employ equal-weight fusion to aggregate behavior-specific embedding. This equal-weight fusion ensures that all behavior types contribute equally to the final representation, while the purification through RRM and ORM ensures that only semantically aligned and optimizational stable signals are preserved. Using these fused embeddings, we compute the final preference score for target behavior prediction: $s(u,i)=\mathbf{z}_{u}^{\top}\mathbf{z}_{i}$.
@@ 4.2. Overall Performance
  • Multi-behavior modeling consistently outperforms single-behavior methods. LightGCN and SimGCL, which rely solely on the target behavior, achieve noticeably lower accuracy compared to multi-behavior models. This validates the effectiveness of incorporating auxiliary behavioral signals, which provide complementary supervision and alleviate target sparsity.
  • Stable and consistent performance across diverse datasets. The performance trends of RMBRec remain stable under different behavioral characteristics. On Taobao, where auxiliary behaviors are moderately noisy, RMBRec achieves substantial gains ( +29.5% on NDCG@10). On TMall, where behaviors are weakly correlated with the target, it maintains competitive performance without degradation. Even on the clean and well-structured Beibei, it continues to outperform all baselines. These results highlight the strong generalization ability of  RMBRec and its robustness to varying behavior quality and distributional patterns.
4.3. Ablation Study
@@
We further investigate the impact of different invariance regularization strategies for the ORM module. Empirical results (detailed in Section C.3) demonstrate that Risk Extrapolation (REx) (Krueger et al., 2021) consistently outperforms gradient-based IRM variants (IRM-V1 and IRM-V2) by enforcing risk consistency across heterogeneous behaviors. Consequently, we adopt REx to ensure global optimization stability in  RMBRec.
@@ Acknowledgements.
This research is supported by the New Cornerstone Science Foundation through the XPLORER PRIZE and the National Natural Science Foundation of China (No. 62272254, No. U25A20445, No. 72188101).
@@ A.2. Detailed Description of  RMBRec Algorithm
Algorithm 1 outlines the complete training procedure of  RMBRec, which systematically integrates multi-behavior learning with dual-level robustness enhancement.
[…]
@@ C.2.4. REx
C.3. Impact of Invariance Regularization Variants
To identify the most effective implementation for our Optimization Robustness Module (ORM), we examine three widely used variants of Invariant Risk Minimization (IRM) (Arjovsky et al., 2019): IRM-V1 (Arjovsky et al., 2019), IRM-V2 (Ahuja et al., 2020), and IRM-REx (Krueger et al., 2021). The comparison results are reported in Table 5.
Variant Taobao TMall Beibei
HR@10 NDCG@10 HR@10 NDCG@10 HR@10 NDCG@10
w/o ORM 0.1321 0.0887 0.1360 0.0766 0.0871 0.0461
IRM-V1 0.1456 0.0962 0.1411 0.0794 0.0913 0.0483
IRM-V2 0.1389 0.0938 0.1304 0.0729 0.0902 0.0478
IRM-REx 0.1462 0.0969 0.1549 0.0838 0.0924 0.0489
As shown in Table 5, all IRM variants substantially outperform the base model without ORM, verifying that invariance regularization effectively suppresses behavior-specific noise. Among them, IRM-REx achieves the best overall performance. This advantage stems from its variance-based penalty, which enforces consistent risks across behavioral environments more stably than the gradient-based penalties in IRM-V1. In contrast, IRM-V2 performs the weakest due to its fixed classifier ($w=1$), which limits the flexibility of invariant representation learning. Based on these findings, we adopt REx as the core instantiation of ORM in  RMBRec.